OpenStudy (anonymous):

Evaluate lim h->0 (1/3+h - 1/3)/h

OpenStudy (anonymous):

I got 1 but is it correct?

OpenStudy (badhi):

If it is, \[\lim \limits_{h \to 0} \frac{\frac{1}{3}+h-\frac{1}{3}}{h}\] yes the answer is 1

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (anonymous):

wait hold no that not the equation :/ it is lim->0 1 over 3 + h then - 1/3 over h sorry

OpenStudy (anonymous):

so would it be -1/9 then?

OpenStudy (badhi):

\[\lim \limits_{h \to 0} \frac{\frac{1}{3+h}-\frac{1}{3}}{h}=\lim \limits_{h \to 0} \frac{-h}{3(3+h)h}=\lim \limits_{h \to 0} \frac{-1}{3(3+h)}=\frac{-1}{9}\] Yes you're correct

OpenStudy (anonymous):

okay thanks :)